OpenStudy (anonymous):

(2-((x+6)/x))/(8+((x-6)/(x+6))) The answer is (x-6)(x+6)/3x(3x+14) I need to know how to get to this answer please help!

OpenStudy (mosaic):

Numerator is: (2x - x - 6) / x = (x-6) / x Denominator is: ( 8(x+6) + (x-6) ) / (x+6) = (8x+48+x-6)/(x+6) = (9x+42)/(x+6) = 3(3x+14)/(x+6) Numerator / denominator = (x-6) / x * (x+6)/{3(3x+14)} = {(x-6)(x+6)} / {3x(3x+14)}

OpenStudy (anonymous):

mosaic thank you sooo much!

OpenStudy (mosaic):

you are welcome.